# Uncertainty-Aware Adaptation of Large Language Models for Protein-Protein Interaction Analysis

### v1

@@ V Conclusion and Discussion
In this study, we presented a novel uncertainty-aware adaptation of LLMs approach for predicting protein-protein interactions across multiple disease contexts. Leveraging fine-tuned LLaMA-3 and BioMedGPT models with LoRA ensemble and Bayesian LoRA, our approach consistently improved prediction accuracy, reliability, and robustness, as confirmed by comprehensive metrics such as negative log-likelihood and calibration error. LoRA ensembles excelled at achieving higher accuracy and reliable uncertainty estimates, while Bayesian LoRA provided well-calibrated predictions. Together, they demonstrated robustness in neurodegenerative, metabolic, and cancer-related PPI tasks. These findings underscore the benefits of incorporating principled uncertainty quantification into parameter-efficient fine-tuning for LLMs. Future work will explore more advanced LLM uncertainty quantification methods and apply this methodology to broader biomedical applications. Potential directions include elucidating disease mechanisms by predicting disrupted protein interactions and, consequently, advance computational protein target discovery for therapeutic design.
@@ V-A Robustness & Predictive Uncertainty Evaluation Metrics
To assess model robustness and predictive uncertainty, we use Negative Log-Likelihood (NLL) and Expected Calibration Error (ECE). NLL evaluates how confidently a model predicts the correct labels. Given a test dataset $\{\bm{x}_{i},y_{i}\}_{i=1}^{N}$, NLL is computed as:
$\text{NLL}=\frac{1}{N}\sum_{i=1}^{N}-\log{\rm P}_{\bm{\theta}}(y_{n}).$
@@
$\text{ECE}=\sum_{m=1}^{M}\frac{|B_{m}|}{n}\left|\text{acc}(B_{m})-\text{conf}(B_{m})\right|.$

### v2

@@ V Conclusion and Discussion
In this study, we presented a novel uncertainty-aware adaptation of LLMs approach for predicting protein-protein interactions across multiple disease contexts. Leveraging fine-tuned LLaMA-3 and BioMedGPT models with LoRA ensemble and Bayesian LoRA, our approach consistently improved prediction accuracy, reliability, and robustness, as confirmed by comprehensive metrics such as negative log-likelihood and calibration error. LoRA ensembles excelled at achieving higher accuracy and reliable uncertainty estimates, while Bayesian LoRA provided well-calibrated predictions. Together, they demonstrated robustness in neurodegenerative, metabolic, and cancer-related PPI tasks. These findings underscore the benefits of incorporating principled uncertainty quantification into parameter-efficient fine-tuning for LLMs. To this end, our method can enable efficient UQ for large-scale AI models in diverse biological tasks, such as agentic AI workflows whose importance is growing in biological data analysis and drug discovery. Furthermore, the quantified uncertainty via our method could be used to assess the reliability of LLM outputs and optimize LLM prompts [56], not only to improve the outputs but also their reliability.
Future work will examine more advanced LLM uncertainty-quantification techniques and extend this framework to broader biomedical applications. Potential directions include leveraging our approach to uncover disrupted protein interactions in disease pathways and, consequently, advance computational protein target discovery for therapeutic design.
@@ V-A Robustness & Predictive Uncertainty Evaluation Metrics
To assess model robustness and predictive uncertainty, we use Negative Log-Likelihood (NLL) and Expected Calibration Error (ECE). NLL evaluates how confidently a model predicts the correct labels. Given a test dataset $\{\bm{x}_{i},y_{i}\}_{i=1}^{N}$, NLL is:
$\text{NLL}=\frac{1}{N}\sum_{i=1}^{N}-\log{\rm P}_{\bm{\theta}}(y_{n}).\vskip-3.00003pt$
@@
$\text{ECE}=\sum_{m=1}^{M}\frac{|B_{m}|}{n}\left|\text{acc}(B_{m})-\text{conf}(B_{m})\right|.\vskip-5.0pt$
@@
V-B Effect of LoRA Rank
In Table II, we present results of the hyperparameter search over the LoRA rank $r\in\{8,16,32\}$ for single LoRA, LoRA ensemble, and Bayesian LoRA across each LLM backbone and PPI dataset combination. For a fair comparison, we use the same selected rank $r$ across all three LoRA-based methods within each LLM backbone–PPI dataset combination. For both LLaMA-3 and BioMedGPT backbones, LoRA-based models achieve overall best performance on the ND-PPI dataset at $r=32$ and on the C-PPI dataset at $r=16$. In contrast, for the M-PPI dataset, LLaMA-3 and BioMedGPT backbones achieve overall best performance across LoRA-based models at $r=16$ and $r=8$, respectively.
ACKNOWLEDGMENT
This research was funded by the Biological and Environmental Research program in the US DOE’s Office of Science under project B&R# KP1601017 and FWP#CC140.